# FourierKAN-GCF: Fourier Kolmogorov-Arnold Network - An Effective and Efficient Feature Transformation for Graph Collaborative Filtering

### v2

@@ 3.2. Fourier KAN
However, KAN is more difficult to train than MLP due to the spline function, which does not satisfy our motivation to provide a more effective and efficient method to replace $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$ part. The essence of KAN is to form an arbitrary function by superposition of multiple nonlinear functions. Our goal can be converted into finding the split from a complex function into multiple relatively simple nonlinear functions. Naturally, the Fourier Coefficients (Nussbaumer and Nussbaumer, 1982) is an interesting choice. Therefore, we propose the following equation:
@@
where $d$ is the dimension number of features. Fourier coefficients $a_{ik}$ and $b_{ik}$ are trainable. Hyper-parameter $g$ is the gridsize, which plays a critical role in controlling the number of terms (frequencies) used in the Fourier series expansion. Specifically, $g$ determines how many different sine and cosine terms are included in the Fourier Coefficients corresponding to each input dimension. The Fourier Coefficients has a significant advantage in computational efficiency and solves the training difficulty caused by the spline function.
@@ 3.3. FourierKAN-GCF
The message passing in our FourierKAN-GCF is defined as:
@@
where $\phi_{F}(\cdot)$ is simplified single layer Fourier KAN function. We remove the unnecessary transform matrix $\mathbf{W}_{1}$ in NGCF and utilize our Fourier KAN function to replace the transform matrix $\mathbf{W}_{2}$. The final user embedding and item embedding are calculated as:
(10) $\mathbf{\hat{e}}_{u}=\mathbf{e}_{u}^{(1)}||\mathbf{e}_{u}^{(2)}||...||\mathbf{e}_{u}^{(L)},\quad\mathbf{\hat{e}}_{i}=\mathbf{e}_{i}^{(1)}||\mathbf{e}_{i}^{(2)}||...||\mathbf{e}_{i}^{(L)},$
where ||||| | is the concatenation operation.
@@ 3.4. Dropout Strategies
To further prevent the over-fitting problem for FourierKAN-GCF, we adopt message dropout and node dropout strategies. For the message dropout, we randomly change some message passing being propagated in Eq 9 by zero, with a probability $1-p_{m}$, where $p_{m}$ is the message dropout ratio. For the node dropout, we randomly drop $1-p_{n}$ nodes of the matrix, where $p_{n}$ is the node dropout ratio.
Table 2. Performance comparison of baselines and FourierKAN-GCF in terms of R@K and N@K. The superscript indicates the improvement is statistically significant where the p-value is less than 0.05.
Datasets 1MOOC 2Games
Metrics R@10 R@20 R@50 N@10 N@20 N@50 R@10 R@20 R@50 N@10 N@20 N@50
BPR-MF 0.2401 0.3353 0.4813 0.1571 0.1898 0.2261 0.0210 0.0369 0.0699 0.0135 0.0183 0.0265
BUIR 0.2354 0.3196 0.4801 0.1589 0.1835 0.2224 0.0227 0.0384 0.0749 0.0143 0.0192 0.0282
NGCF 0.2453 0.3361 0.4799 0.1722 0.1894 0.2349 0.0231 0.0379 0.0782 0.0142 0.0196 0.0274
LR-GCCF 0.2466 0.3336 0.4809 0.1678 0.1938 0.2294 0.0253 0.0440 0.0815 0.0171 0.0224 0.0317
LightGCN 0.2402 0.3307 0.4773 0.1581 0.1811 0.2217 0.0271 0.0447 0.0844 0.0174 0.0227 0.0326
UltraGCN 0.2372 0.3194 0.4701 0.1737 0.1962 0.2307 0.0277 0.0459 0.0844 0.0173 0.0230 0.0331
IMP-GCN 0.2170 0.2788 0.4183 0.1531 0.1717 0.2057 0.0273 0.0461 0.0839 0.0171 0.0232 0.0323
w/o MD 0.2518 0.3523 0.4912 0.1814 0.2116 0.2449 0.0267 0.0452 0.0825 0.0165 0.0221 0.0314
w/o ND 0.2603 0.3527 0.4839 0.1800 0.2071 0.2439 0.0257 0.0452 0.0809 0.0161 0.0219 0.0309
KAN-GCF 0.2536 0.3417 0.4984 0.1758 0.2024 0.2396 0.0272 0.0451 0.0837 0.0172 0.0229 0.0325
FourierKAN-GCF 0.2595 0.3564 0.5065 0.1866 0.2147 0.2462 0.0287 0.0473 0.0856 0.0176 0.0252 0.0342
1Dataset can be accessed at http://moocdata.cn. 2Dataset can be accessed at http://jmcauley.ucsd.edu/data/amazon/links.html.
@@ 3.5. Model Training
To learn model parameters, we adopt the Bayesian Personalized Ranking (BPR) (Rendle et al., 2012) loss function as our optimization criterion. The core objective of BPR is to enhance the divergence in the predictive preference between positive and negative items within each user-item triplet $(u,i_{p},i_{n})\in\mathcal{D}$, where $\mathcal{D}$ signifies the collection of training data, the term positive item $p$ pertains to an item with which the user $u$ has interacted, whereas the negative item $n$ is selected randomly from the pool of items without interaction with user $u$.
@@
MOOC 82535 1302 458453 99.57%
Games 50677 16897 454529 99.95%
4. Experiment
@@ 4.1. Datasets and Evaluation Metrics
We conduct our experiments on two real-world datasets: MOOC and Anazon Video Games (Games for short). Details can be found in Table 3. For a fair comparison, we follow previous work (Zhou et al., 2023b) to sort all the observed user-item interactions of each dataset in chronological order based on the interaction timestamps. Then, we split each dataset with a ratio of 7:1:2 for training, validation, and testing. Regarding evaluation metrics, we adopt two well-established metrics: Recall@K (R@K) and NDCG@K (N@K). We report the average metrics of all users in the test dataset among $K$ = 10, $K$ = 20, and $K$ = 50.
@@ 4.2. Baselines and Experimental Settings
To verify the effectiveness of FourierKAN-GCF, we select two traditional matrix factorization models (BPR-MF (Rendle et al., 2012) and BUIR (Lee et al., 2021)) and five state-of-the-art GCN-based models (NGCF (Wang et al., 2019), LR-GCCF (Chen et al., 2020), LightGCN (He et al., 2020), UltraGCN (Mao et al., 2021), and IMP-GCN (Liu et al., 2021)) for comparison. For a fair comparison, we implement all baselines under a unified framework (Zhou et al., 2023b), fix the embedding size of both users and items to 64 for all models, initialize the embedding parameters with the Xavier initialization (Glorot and Bengio, 2010), and use Adam (Kingma and Ba, 2014) as the optimizer. Besides, we carefully tune the hyper-parameters of each baseline following their published papers to find optimal settings. For our FourierKAN-GCF, we tune the $\lambda$ of $L_{2}$ regularization term in {0, 1e-2, 1e-1, 1, 10}, and tune the layer number $L$ from 1 to 4. The gridsize $g$ of Fourier KAN is searched from {1, 2, 4, 8}. We further grid search the message dropout ratio $p_{m}$ and node dropout ratio $p_{n}$ in {0.0, 0.1, 0.2, 0.3}.
@@ 4.3. Performance Comparison
The results of our experiments are listed in Table 2. We can identify that the majority of GCN-based methods outperform traditional matrix factorization methods, which is a common trend. Our FourierKAN-GCF outperforms all baselines in two datasets across various metrics. We owe our superiority to FourierKAN, which is easier to train and has greater representation power than MLP. It is worth noting that FourierKAN-GCF can adjust the training difficulty by adjusting the grid size $g$.
@@ 4.5. Sensitivity Analysis
To analyze the hyper-parameter sensitivity of FourierKAN-GCF, we test the performance of FourierKAN-GCF on two datasets with different hyper-parameters. The optimal layer number $L$ is 3 on all datasets, shown in Figure LABEL:fig:l. Besides, Figure LABEL:fig:g demonstrates that for MOOC and Games datasets, optimal ($p_{m}$, $p_{n}$) pairs are (0.2, 0.2) and (0.1, 0.1), respectively. As Figure LABEL:fig:dropout illustrated, for the relatively dense dataset MOOC, $g$ = 4 is the best grid size. For the relatively sparse dataset Games, $g$ = 2 is the best grid size. This further demonstrates that FourierKAN can be easily adapted to various datasets.
@@ 5. Conclusion
In this paper, we revisit the processes of feature transformation and nonlinear operations during message passing in the NGCF. We assert that feature transformation in NGCF can enhance the interaction representation and boost the performance of GCN but add complexity in training. To reduce the training difficulty while benefiting from these representational enhancements, we introduce a novel feature transformation method called FourierKAN. Inspired by KAN, FourierKAN employs a Fourier Coefficients instead of the Spline function in standard KAN, which further reduces the difficulty of training while increasing the representation power. Furthermore, based on FourierKAN, we introduce a simple and effective graph-based recommendation model (FourierKAN-GCF) with extra dropout strategies. Extensive experiments on public datasets demonstrate the superiority of our model over the state-of-the-art methods.

### v3

@@ 3.2. Fourier KAN
To further reduce the training difficulty and be able to adopt to different scenarios. Our goal can be converted into finding the split from a complex function into multiple relatively simple nonlinear functions. Naturally, the Fourier Coefficients (Nussbaumer and Nussbaumer, 1982) is a potential choice. Therefore, we propose the following equation:
@@
where $d$ is the dimension number of features. Fourier coefficients $a_{ik}$ and $b_{ik}$ are trainable. Hyper-parameter $g$ is the gridsize, which plays a critical role in controlling the number of terms (frequencies) used in the Fourier series expansion. Specifically, $g$ determines how many different sine and cosine terms are included in the Fourier Coefficients corresponding to each input dimension. The Fourier Coefficients has a significant advantage in computational efficiency and reduces the training difficulty caused by the spline function.
@@ 3.3. FourierKAN-GCF
The message passing in FourierKAN-GCF is defined as:
@@
where $\phi_{F}(\cdot)$ is simplified single layer Fourier KAN function. We remove the unnecessary transform matrix $\mathbf{W}_{1}$ in NGCF and utilize our Fourier KAN function to replace the transform matrix $\mathbf{W}_{2}$. The final user embedding and item embedding are calculated by $\mathbf{\hat{e}}_{u}=\{\mathbf{e}_{u}^{(1)}||\mathbf{e}_{u}^{(2)}||...||\mathbf{e}_{u}^{(L)}\}$ and $\mathbf{\hat{e}}_{i}=\{\mathbf{e}_{i}^{(1)}||\mathbf{e}_{i}^{(2)}||...||\mathbf{e}_{0}^{(L)}\}$, where $||$ is the concatenation operation.
@@ 3.4. Dropout Strategies
To mitigate overfitting in FourierKAN-GCF, we employ message dropout and node dropout strategies, similar to NGCF (Wang et al., 2019). In message dropout, a fraction $1-p_{m}$ of message passing in Eq. 5 is randomly set to zero, where $p_{m}$ is the dropout ratio. In node dropout, $1-p_{n}$ of the nodes in the matrix are randomly dropped, with $p_{n}$ as the dropout ratio.
@@ 3.5. Model Training
For model optimization, we adopt the Bayesian Personalized Ranking (BPR) (Rendle et al., 2009) loss function as our optimization criterion. The core objective of BPR is to enhance the divergence in the predictive preference between positive and negative items within each user-item triplet $(u,i_{p},i_{n})\in\mathcal{D}$, where $\mathcal{D}$ signifies the collection of training data, the term positive item $p$ pertains to an item with which the user $u$ has interacted, whereas the negative item $n$ is selected randomly from the pool of items without interaction with user $u$.
@@
4. Experiment
In this section, we compare the performance of FourierKAN-GCF with popular graph-based backbone models and demonstrate its compatibility with advanced self-supervised graph-based models.
@@ 4. Experiment
MOOC 82,535 1,302 458,453 99.57%
Amazon 50,677 16,897 454,529 99.95%
Gowalla 29,859 40,989 1,027,464 99.92%
@@ 4.1. Datasets and Evaluation Metrics
Table 3. Performance comparison of baselines and FourierKAN-GCF in terms of R@K and N@K. The superscript indicates the improvement is statistically significant where the p-value is less than 0.05. #T denotes seconds per epoch.
Datasets MOOC Amazon Gowalla
Metrics R@20 R@50 N@20 N@50 #T R@20 R@50 N@20 N@50 #T R@20 R@50 N@20 N@50 #T
NGCF 0.3361 0.4799 0.1894 0.2349 5.2s 0.0379 0.0782 0.0196 0.0274 4.4s 0.1755 0.2811 0.1013 0.1270 37.5s
LR-GCCF 0.3336 0.4809 0.1938 0.2294 5.1s 0.0440 0.0815 0.0224 0.0317 4.7s 0.1803 0.2971 0.1101 0.1369 40.1s
LightGCN 0.3307 0.4773 0.1811 0.2217 4.3s 0.0447 0.0844 0.0227 0.0326 3.5s 0.1830 0.3044 0.1152 0.1414 26.4s
UltraGCN 0.3194 0.4701 0.1962 0.2307 4.9s 0.0459 0.0844 0.0230 0.0331 4.0s 0.1798 0.2909 0.1059 0.1328 33.8s
IMP-GCN 0.2788 0.4183 0.1717 0.2057 54.2s 0.0461 0.0839 0.0232 0.0323 44.9s 0.1808 0.2932 0.1060 0.1345 143.6s
KAN-GCF 0.3417 0.4984 0.2024 0.2396 4.9s 0.0451 0.0837 0.0229 0.0325 4.0s 0.1922 0.3023 0.1142 0.1403 32.9s
FourierKAN-GCF 0.3564 0.5065 0.2147 0.2462 4.6s 0.0473 0.0856 0.0252 0.0342 3.7s 0.1962 0.3077 0.1179 0.1436 29.8s
w/o MD 0.3523 0.4912 0.2116 0.2449 4.5s 0.0452 0.0825 0.0221 0.0314 3.7s 0.1920 0.3031 0.1150 0.1408 29.5s
w/o ND 0.3527 0.4839 0.2071 0.2439 4.5s 0.0452 0.0809 0.0219 0.0309 3.6s 0.1908 0.3017 0.1133 0.1389 29.6s
Table 4. Performance comparison of different backbones on advanced recommendation models in terms of R@K and N@K.
Datasets MOOC Amazon Gowalla
Metrics Backbones R@20 R@50 N@20 N@50 R@20 R@50 N@20 N@50 R@20 R@50 N@20 N@50
SimGCL LightGCN 0.3503 0.5032 0.2109 0.2428 0.0462 0.0837 0.0230 0.0327 0.2028 0.3126 0.1184 0.1459
FourierKAN-GCF 0.3639 0.5188 0.2216 0.2543 0.0482 0.0868 0.0258 0.0352 0.2084 0.3163 0.1234 0.1521
LightGCL LightGCN 0.2742 0.4139 0.1697 0.2018 0.0453 0.0818 0.0225 0.0318 0.1888 0.2976 0.1081 0.1348
FourierKAN-GCF 0.2823 0.4291 0.1733 0.2072 0.0479 0.0864 0.0250 0.0347 0.2007 0.3100 0.1198 0.1452
RecDCL LightGCN 0.3531 0.5009 0.2113 0.2423 0.0469 0.0848 0.0235 0.0334 0.1993 0.3052 0.1160 0.1433
FourierKAN-GCF 0.3608 0.5120 0.2175 0.2514 0.0484 0.0863 0.0253 0.0349 0.2051 0.3113 0.1203 0.1472
We conduct experiments on three real-world datasets: MOOC, Amazon Video Games (Amazon), and Gowalla. Details can be found in Table 2. For a fair comparison, we sort all observed user-item interactions chronologically based on the interaction timestamps. Then, we split each dataset with a ratio of 7:1:2 for training, validation, and testing. Regarding evaluation metrics, we adopt two well-established metrics (Järvelin and Kekäläinen, 2002): Recall@K (R@K) and NDCG@K (N@K).
@@ 4.2. Baselines and Experimental Settings
To verify the effectiveness of FourierKAN-GCF, we select five GCN-based backbone models (NGCF (Wang et al., 2019), LR-GCCF (Chen et al., 2020), LightGCN (He et al., 2020), UltraGCN (Mao et al., 2021), and IMP-GCN (Liu et al., 2021)) as baselines. Moreover, FourierKAN-GCF can also be a more powerful backbone to enhance existing advanced self-supervised enhanced graph-based models (SimGCL (Yu et al., 2022), LightGCL (Cai et al., 2023), and RecDCL (Zhang et al., 2024)). For a fair comparison, we fix the embedding size of both users and items to 64 for all models, initialize embedding parameters with the Xavier initialization (Glorot and Bengio, 2010), and use Adam (Kingma and Ba, 2015) as optimizer. Besides, we tune the hyper-parameters of each baseline following their published papers. For FourierKAN-GCF, we fix the $\lambda$ = $1$ for $L_{2}$ regularization, and tune layer number $L$ from 1 to 4. The gridsize $g$ is searched from {1, 2, 4, 8}. Message and node dropout ratios $p_{m}$ and $p_{n}$ is searched from {0.0, 0.1, 0.2, 0.3}.
@@ 4.3. Performance Comparison
The results of our experiments are listed in Table 3. Our FourierKAN-GCF outperforms all baselines in three datasets across various metrics. Moreover, FourierKAN-GCF is second only to LightGCN in efficiency, which demonstrates it effectively reduces the training difficulty associated with feature transformation. We owe our superiority to FourierKAN, which is easier to train and has greater representation power than MLP. Note that FourierKAN-GCF can adjust the training difficulty by adjusting the grid size $g$.
4.4. Compatibility Analysis
We further evaluate the compatibility of FourierKAN-GCF with advanced self-supervised graph-based models by replacing their backbones. As shown in Table 4, adopting FourierKAN-GCF significantly improves performance, highlighting the importance of feature transformation in GCF and establishing FourierKAN-GCF as an effective solution.
@@ 4.5. Ablation Study
(a) Layer number $L$ (Recall@20)
(b) Grid size $g$ (Recall@20)
Figure 2. Effective of layer number $L$ and grid size $g$.
(a) MOOC
(b) Amazon
(c) Gowalla
Figure 3. Study of dropout ratio pair $p_{m}$ and $p_{n}$ (Recall@20).
@@ 4.6. Sensitivity Analysis
To analyze the hyper-parameter sensitivity of FourierKAN-GCF, we test the performance of FourierKAN-GCF on three datasets with different hyper-parameters. The optimal layer number $L$ is 3 on all datasets, shown in Figure 2(a). Besides, Figure 2(b) demonstrates that for MOOC dataset, the optimal ($p_{m}$, $p_{n}$) pair is (0.2, 0.2), and for Amazon and Gowalla datasets, the optimal pairs are all (0.1, 0.1). As Figure 3 illustrated, for the relatively dense dataset MOOC, $g$ = 4 is the best grid size. For the relatively sparse datasets Amazon and Gowalla, $g$ = 2 is the best grid size. This further demonstrates that FourierKAN can be easily adapted to various datasets.
@@ 5. Conclusion
In this paper, we revisit feature transformation and nonlinear operations in the message-passing mechanism of GCNs. While feature transformation enhances interaction representation and boosts performance, it increases training complexity. To this end, we introduce a new feature transformation for GCF called FourierKAN. Inspired by KAN, FourierKAN employs the Fourier Coefficients instead of the Spline function in standard KAN. We further propose a simple yet effective GCF model (FourierKAN-GCF), which reduces the difficulty of training. In addition, FourierKAN-GCF can be integrated into existing advanced self-supervised models as a backbone, replacing their original backbone to achieve enhanced performance. Extensive experiments on public datasets verify the superiority of our model over the advanced methods.
GenAI Usage Disclosure
No GenAI tools were used in any stage of the research, nor in the writing.
Acknowledgements.
This work was supported by the Hong Kong UGC General Research Fund no. 17203320 and 17209822, and the project grants from the HKU-SCF FinTech Academy.